The average price of a firms stock is given in the table below for each of the last two years, together with the firms earnings in each of those years. Using the average of the P/E multiples for the last 2 years as the relevant P/E benchmark, determine an average expected stock price for the coming year (year 3) if the firms earnings are expected to grow by 8.4% from those in year 2.

Make sure to perform intermediate calculations to 3 decimal places, and record your final answer to two decimal places

Year 1 Year 2
Average Stock Price 24.8 27.4
Earnings Per share (EPS) 1.55 1.86
Answer #1

P/E year 1 =Stock Price/EPS =24.80/1.55=16
P/E year 2 =Stock Price/EPS =27.40/1.86=14.7312
EPS year 3=EPS year 2*(1+g) =1.86*(1+8.4%) =2.016
Average P/E =(16+14.7312)/2=15.366
Average Stock Price =(EPS year 3*P/E average) =2.016*15.366 =30.98